OpenStudy (anonymous):

if the sum of 5 consecutive positive integers is w, in terms of w, which if the following represents the sum of the next 5 consecutive positive integers? a) w+5 b) 5w+5 c) 5w+25 d) w+25 e)w^2 +25

OpenStudy (anonymous):

i tried it but dont got the answer.... :(

OpenStudy (anonymous):

@DoYourHomework ?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

Take this as an example \[1+2+3+4+5=15\] Let w = 15 The sum of the next 5 consecutive numbers is \[6+7+8+9+10=40\] In terms of w, which is equal to 15. 40 equals w+25 = 40

OpenStudy (anonymous):

Does it make sense?

OpenStudy (anonymous):

oh... thanx lol my retardedness made me get 5w+15